#opnfv-meeting: BGS daily release readiness status meeting

Meeting started by frankbrockners at 15:59:34 UTC (full logs).

Meeting summary

    1. Frank Brockners (frankbrockners, 16:00:02)
    2. pbandzi (pbandzi, 16:00:17)
    3. Jonas Bjurel ([1]JonasB, 16:01:04)
    4. Heather Kirksey (HKirksey, 16:01:49)
    5. Bryan Sullivan (bryan_att, 16:03:35)
    6. Tim Rozet (trozet, 16:03:39)

  1. Updates on BGS status - updating https://wiki.opnfv.org/releases/arno/releasereadiness (frankbrockners, 16:05:23)
    1. workaround added to deploy so that vbox VM is now working on jumphost. Still not sure what the root cause is, but we suspect it is UCS adding a VLAN tag egress to the host (trozet, 16:06:42)
    2. VLAN header with VID 0 from UCS is due to no VLAN being configured from UCS-> Fabric Interconnect, Frank to talk about changes to the LF setup (trozet, 16:07:18)
    3. IPMI now working after LF changes to the firewall, and verified Foreman can control the nodes (trozet, 16:07:39)
    4. testing out building a node right now, will try to deploy nodes today (trozet, 16:08:00)
    5. each installer should cleanup the kernel mods it is using after a job, so that hypervisors dont conflict. Example: kvm modules colliding with vboxdrv (trozet, 16:12:43)
    6. documentation is 100% ready for review in git ([1]JonasB, 16:14:13)
    7. Were finalizing the HW adapter for Eri lab and hope to autodeploy on Mon-Tue ([1]JonasB, 16:14:52)
    8. Started to work on HW adapter for LF-lab ([1]JonasB, 16:15:17)
    9. Centos support in, a small issue with HA for centos, troubleshooting ongoing. ([1]JonasB, 16:16:05)
    10. : Jose is working with Tempest tests for Fuel), should be 100% reusable for Foreman ([1]JonasB, 16:16:52)

  2. Testing for BGS (frankbrockners, 16:17:52)
    1. Current understanding is: Functest will supply a set of tests (Rally/Tempest and Robot). Functest will also take care of installing Robot and Tempest/Rally on the jumphosts of POD1 and POD2. (frankbrockners, 16:19:00)
    2. BGS is not immediately involved with testing - only as that we have to make sure their tests exec successfully (frankbrockners, 16:19:44)
    3. rally & robot have been installed on jhs (per fdegir) (frankbrockners, 16:20:23)

  3. Updates to LF HW config (frankbrockners, 16:23:59)
    1. One thing related to LF. The admin network will need to be routable to the public network so that the PXE nodes can get out to the internet. (trozet, 16:25:03)
    2. For proper operation, there is a need to have all traffic tagged *within* the fabric (you won't notice on the vNIC). Right now, we don't do this in all cases. (frankbrockners, 16:25:05)
    3. E.g. ...9.64/26 isn't mapped to a VLAN in the fabric. (frankbrockners, 16:26:17)
    4. In addition, there is a need to fully isolate PODs (so that we don't confuse each other) - that would mean subnetting the current .8.0/24 and .9.0/24 further. (frankbrockners, 16:27:10)
    5. Recommendation from my end: We gather all re-config requirements on a wiki or etherpad and once complete - we do the change in a single maintenance window - because it will impact connectivity to the jumphosts etc. (frankbrockners, 16:28:46)
    6. HW config changes required to be gathered in a section on the wiki https://wiki.opnfv.org/get_started/lflab_hosting - once stable, we'll move the entire page to git (needed for Arno anyway) (frankbrockners, 16:36:53)

Meeting ended at 16:38:06 UTC (full logs).

Action items

  1. (none)

People present (lines said)

  1. frankbrockners (40)
  2. [1]JonasB (21)
  3. trozet (18)
  4. collabot (3)
  5. bryan_att (1)
  6. pbandzi (1)
  7. fdegir (1)
  8. HKirksey (1)

Generated by MeetBot 0.1.4.